OpenStudy (anonymous):

Write the polar form of the complex number 3-2i, using θ such that 0 ≤ θ ≤ 360º. Round numerical values to the nearest hundredth.

OpenStudy (anonymous):

you need two numbers to find this and put it in the form \[3-2i=r(\cos(\theta)+i\sin(\theta))\] namely \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r\) is straightforward, it is \[r=\sqrt{a^2+b^2}\] in your case \(a=3,b=2\)

OpenStudy (anonymous):

well actually \(b=-2\) but it makes no difference when you square it what do you get for \(r\) ?

OpenStudy (anonymous):

2.2? i think i did it wrong though..

OpenStudy (anonymous):

yeah i think so you need \(r=\sqrt{3^2+2^2}\)

OpenStudy (anonymous):

3.6

OpenStudy (anonymous):

i guess, it is actually \(\sqrt{13}\)

OpenStudy (anonymous):

if you round to the nearest 100th it is \(3.61\)

OpenStudy (anonymous):

now for \(\theta\) since \(3-2i\) is in quadrant 4, you can use \[\theta=\tan^{-1}(\frac{b}{a})=\tan^{-1}(-\frac{2}{3})\]

OpenStudy (anonymous):

and of course a calculator to find it

OpenStudy (anonymous):

-33.69

OpenStudy (anonymous):

maybe if you are working in degrees usually these are just numbers (radians) but if you are working in degrees then you are right

OpenStudy (anonymous):

you get a final answer of \[3-2i=3.61(\cos(-33.69)+i\sin(-33.69))\]

OpenStudy (anonymous):

thank you!